# Geo-Registration of Terrestrial LiDAR Point Clouds with Satellite Images without GNSS

### v3

@@ Abstract
Accurate geo-registration of LiDAR point clouds remains a significant challenge in urban environments where Global Navigation Satellite System (GNSS) signals are denied or degraded. Existing methods typically rely on real-time GNSS and Inertial Measurement Unit (IMU) data, which require pre-calibration and assume stable signals. However, this assumption often fails in dense cities, resulting in localization errors. To address this, we propose a structured geo-registration method that accurately aligns LiDAR point clouds with satellite images, enabling frame-wise geo-registration and city-scale 3D reconstruction without prior localization. Our method uses a pre-trained Point Transformer to segment road points, then extracts road skeletons and intersections from the point cloud and the satellite image. Global alignment is achieved through rigid transformation using corresponding intersection points, followed by local non-rigid refinement with radial basis function (RBF) interpolation. Elevation discrepancies are corrected using terrain data from the Shuttle Radar Topography Mission (SRTM). To evaluate geo-registration accuracy, we measure the absolute distances between the roads extracted from the two modalities. Our method is validated on the KITTI benchmark and a newly collected dataset of Perth, Western Australia. On KITTI, our method achieves a mean planimetric alignment error of 0.69m, representing 50% improvement over the raw KITTI data. On Perth dataset, it achieves a mean planimetric error of 2.17m from GNSS values extracted from Google Maps, corresponding to 57.4% improvement over rigid alignment. Elevation correlation improved by 30.5% (KITTI) and 55.8% (Perth). A demonstration video is available at: Demo Video.
@@ I Introduction
To address these problems, this paper introduces a structured Geo-Registration and Spatial Correction method based on topology-driven skeleton extraction, spatial transformations, and elevation correction. The proposed method begins with semantic segmentation to extract road points from raw LiDAR point clouds, which are then projected onto a plane to generate a simplified road skeleton. In parallel, high-resolution satellite maps are processed to produce matching road skeletons and intersection keypoints. Alignment is performed hierarchically: first, a rigid transformation aligns the two modalities using matched intersections, ensuring global consistency; then, a non-rigid refinement using radial basis function (RBF) interpolation corrects local distortions. Finally, terrain-aware elevation correction aligns the point cloud vertically using Shuttle Radar Topography Mission (SRTM) data, achieving full 3D spatial consistency. To comprehensively assess registration performance, we introduce two metrics that evaluate global consistency and local precision. Experiments are performed on the KITTI benchmark [39] and a newly collected dataset in Perth city, Western Australia. On KITTI, our method reduces the mean planimetric alignment error from 1.38m to 0.69m (50% improvement) and on Perth dataset, it achieves 2.17m against GNSS values extracted from Google Maps (57.4% improvement). Elevation correlation improved by 30.5% on KITTI and 50.4% on the Perth dataset. A Demo Video of the geo-registered 3D Perth city map shows our qualitative results.
@@
    We introduce a novel problem formulation of city-scale geo-registeration of terrestrial LiDAR point clouds without GNSS data, addressing practical scenarios where accurate satellite positioning is unavailable and data re-acquisition is not feasible. We will release the geo-registered Perth data to facilitate future research.
@@ II Literature Review
Current approaches for GNSS coordinate annotation of point clouds primarily rely on multi-sensor fusion techniques that integrate GNSS with LiDAR and IMU data [11, 23] during collection. However, these methods face limitations in urban environments due to the "urban canyon effect” [38], where signal occlusion degrades GNSS accuracy [18]. Alternative approaches include 3D map-based global localization (MGL) using pre-existing 3D maps [37], which requires pre-constructed 3D point cloud maps for the corresponding mapping area. Additionally, SLAM-based methods that operate independently in GNSS-denied environments [15, 16] have been developed. While these techniques can provide relative positioning accuracy, they struggle with absolute georeferencing without external reference constraints.
@@
Rigid alignment corrects global misalignments but cannot address local geometric distortions inherent to mobile LiDAR mapping. Radial basis function (RBF) methods, such as thin-plate splines [50, 51], enable continuous non-rigid warping guided by control points, and have been applied in limited urban contexts [52, 53]. The selection of topological control points, such as road intersections, is critical to preserving spatial coherence during non-rigid correction.
@@
Current methods face persistent limitations: reliance on GNSS or manual intervention, sensitivity to urban occlusion and topological abstraction, and inadequate handling of local distortions in 3D–2D registration. There is a clear need for a fully automatic, topology-driven framework that achieves robust urban geo-registration and correction using only LiDAR and map data, and can be rigorously evaluated on urban-scale datasets. Our work addresses these gaps with a unified approach combining semantic segmentation, skeleton-based topology matching, hierarchical transformation, and standardized quantitative assessment.
@@ III Methodology
Adaptive voxel downsampling. To reduce computational cost while preserving geometry, the segmented road cloud is adaptively voxelized. The voxel size is defined as $s=\alpha\cdot\left(\frac{\mathrm{vol}(P)}{|P|}\right)^{1/3}$, where $\mathrm{vol}(P)$ is the bounding-box volume, $|P|$ the point count, and $\alpha$ a scaling factor. One representative point per voxel is retained, substantially reducing density while maintaining structural fidelity.
@@ III-C Rigid alignment based on feature extraction
$\begin{split}\theta=-\Big[&\arctan 2(p_{k}^{1}[1]-p_{i}^{1}[1],p_{k}^{1}[0]-p_{i}^{1}[0])\\ &-\arctan 2(p_{l}^{2}[1]-p_{j}^{2}[1],p_{l}^{2}[0]-p_{j}^{2}[0])\Big]\end{split}$ (3)
@@ III-D Non-rigid transformation using radial basis functions
where $\mathbf{w}_{i}$ are weights determined by the displacements $\mathbf{t}_{i}-\mathbf{s}_{i}$, $\phi(\cdot)$ is the radial basis kernel, and $\mathbf{P}(\mathbf{x})$ is a polynomial term for global affine adjustment. We adopt the thin-plate spline (TPS) kernel, defined as $\phi(r,\epsilon)=(r^{2}+\epsilon^{2})^{1/2}$, with the regularization parameter $\epsilon=0.6$.
@@ IV-B Experimental Setup
All experiments were performed on a workstation with Intel Core i9-10900K CPU (3.7 GHz), 32 GB DDR4 memory, and an NVIDIA RTX 3080 GPU (12 GB VRAM). Implementation was done in Python 3.9 using Anaconda environment. Core libraries included NumPy, SciPy, Open3D, scikit-learn, Matplotlib, and OpenCV 4.5. Deep learning inference for semantic segmentation was based on PyTorch 1.13 with CUDA 11.6 support. Algorithm parameters were chosen as follows: keypoint matching threshold = 8 pixels, skeleton extraction density = 40, RBF interpolation parameter $\epsilon=5$, and correspondence distance threshold = 30 pixels. Point cloud preprocessing used 20 nearest neighbors and a std multiplier of 2.0 for outlier removal.
@@ V Conclusion
We presented a method for geo-registering terrestrial LiDAR point clouds, captured without GNSS, to satellite images or maps. Our method is also useful for correcting GNSS errors as demonstrated on the KITTI dataset. Adhering to a robust pipeline, with meticulous attention to detail and precision, our method ensures highly accurate alignment of road structures extracted from the point clouds and the satellite view as demonstrated on two datasets. It also integrates elevation data to correct vertical errors in the geo-registered 3D point cloud. With road structures as the main alignment features, our method is most suitable for autonomous driving applications and robust to changes in the buildings. Finally, we provide geo-referencing data for the publicly available complete Perth CBD dataset and the five longest sequences of the KITTI dataset to facilitate future research in this direction.

### v4

@@ Abstract
Accurate geo-registration of LiDAR point clouds remains a significant challenge in urban environments where Global Navigation Satellite System (GNSS) signals are denied or degraded. Existing methods typically rely on real-time GNSS and Inertial Measurement Unit (IMU) data, which require pre-calibration and assume stable signals. However, this assumption often fails in dense cities, resulting in localization errors. To address this, we propose a structured post-hoc geo-registration method that accurately aligns LiDAR point clouds with satellite images. The proposed approach targets point cloud datasets where reliable GNSS information is unavailable or degraded, enabling city-scale geo-registration as a post-processing solution. Our method uses a pre-trained Point Transformer to segment road points, then extracts road skeletons and intersections from the point cloud and the satellite image. Global alignment is achieved through rigid transformation using corresponding intersection points, followed by local non-rigid refinement with radial basis function (RBF) interpolation. Elevation discrepancies are corrected using terrain data from the Shuttle Radar Topography Mission (SRTM). To evaluate geo-registration accuracy, we measure the absolute distances between the roads extracted from the two modalities. Our method is validated on the KITTI benchmark and a newly collected dataset of Perth, Western Australia. On KITTI, our method achieves a mean planimetric alignment error of 0.69 m, corresponding to a 50% reduction in global geo-registration bias compared to the raw KITTI annotations. On Perth dataset, it achieves a mean planimetric error of 2.17m from GNSS values extracted from Google Maps, corresponding to 57.4% improvement over rigid alignment. Elevation correlation factor improved by 30.5% (KITTI) and 55.8% (Perth). A demonstration video is available at: Demo Video.
@@ I Introduction
To address these problems, this paper introduces a structured post-hoc geo-registration and spatial correction method based on topology-driven skeleton extraction, spatial transformations, and elevation correction, specifically targeting terrestrial LiDAR datasets that lack reliable geographic referencing. The proposed method begins with semantic segmentation to extract road points from raw LiDAR point clouds, which are then projected onto a plane to generate a simplified road skeleton. In parallel, high-resolution satellite maps are processed to produce matching road skeletons and intersection keypoints. Alignment is performed hierarchically: first, a rigid transformation aligns the two modalities using matched intersections, ensuring global consistency; then, a non-rigid refinement using radial basis function (RBF) interpolation corrects local distortions. Finally, terrain-aware elevation correction aligns the point cloud vertically using Shuttle Radar Topography Mission (SRTM) data, achieving full 3D spatial consistency. To comprehensively assess registration performance, we introduce two metrics that evaluate global consistency and local precision. Experiments are performed on the KITTI benchmark [39] and a newly collected dataset in Perth city, Western Australia. On KITTI, our method reduces the mean planimetric alignment error from 1.38m to 0.69m (50% improvement) and on Perth dataset, it achieves 2.17m against GNSS values extracted from Google Maps (57.4% improvement). Elevation correlation improved by 30.5% on KITTI and 50.4% on the Perth dataset. A Demo Video of the geo-registered 3D Perth city map shows our qualitative results.
@@
    We formulate and address a post-hoc city-scale geo-registration problem for terrestrial LiDAR point clouds, targeting practical scenarios where GNSS information is unavailable, unreliable, or erroneous during data acquisition.
@@ II Literature Review
Current approaches for GNSS coordinate annotation of point clouds primarily rely on multi-sensor fusion techniques that integrate GNSS with LiDAR and IMU data [11, 23] during collection. However, these methods face limitations in urban environments due to the "urban canyon effect” [38], where signal occlusion degrades GNSS accuracy [18]. Alternative approaches include 3D map-based global localization (MGL) using pre-existing 3D maps [37], which requires pre-constructed 3D point cloud maps for the corresponding mapping area. Additionally, SLAM-based methods that operate independently in GNSS-denied environments [15, 16] have been developed. While these techniques achieve accurate relative positioning, they do not address the geo-registration of existing point cloud datasets collected without reliable GNSS to global geographic coordinates.
@@
Rigid alignment corrects global misalignments but cannot address local geometric distortions inherent to mobile LiDAR mapping. RBF methods, such as thin-plate splines [50, 51], enable continuous non-rigid warping guided by control points, and have been applied in limited urban contexts [52, 53]. The selection of topological control points, such as road intersections, is critical to preserving spatial coherence during non-rigid correction.
@@
Existing approaches face key limitations for post-hoc geo-registration of terrestrial LiDAR point clouds. GNSS/IMU fusion and map-based global localization typically require reliable positioning information during data acquisition, while SLAM and LIO methods primarily focus on relative pose estimation and do not directly provide global geographic coordinates. Most existing 3D–2D registration approaches further assume coarse initial alignment or ground control points, limiting their applicability in large-scale urban environments. Moreover, rigid alignment alone cannot adequately correct the local distortions commonly observed in large-scale mobile mapping. These limitations highlight the need for a fully automatic framework that enables robust urban geo-registration by leveraging LiDAR data and publicly available satellite imagery, particularly in scenarios where reliable GNSS information is unavailable. Our work addresses these gaps with a unified approach that combines semantic road segmentation, skeleton-based topology matching, hierarchical rigid-to-non-rigid transformation, and terrain-aware elevation correction.
@@ III Methodology
Adaptive voxel downsampling. To reduce computational cost while preserving geometry, the segmented road cloud is adaptively voxelized. The voxel size is defined as $s=\alpha\cdot\left(\frac{\mathrm{vol}(P)}{|P|}\right)^{1/3}$, where $\mathrm{vol}(P)$ is the bounding-box volume, $|P|$ the point count, and $\alpha$ a scaling factor. In all experiments, $\alpha$ is set to 1.1, such that the voxel size is slightly larger than the average point spacing. One representative point per voxel is retained, substantially reducing density while maintaining structural fidelity.
@@ III-C Rigid alignment based on feature extraction
$\begin{split}\theta=-\Big[&\operatorname{atan2}(p_{k}^{1}[1]-p_{i}^{1}[1],\,p_{k}^{1}[0]-p_{i}^{1}[0])\\ &-\operatorname{atan2}(p_{l}^{2}[1]-p_{j}^{2}[1],\,p_{l}^{2}[0]-p_{j}^{2}[0])\Big]\end{split}$ (3)
@@ III-D Non-rigid transformation using radial basis functions
where $\mathbf{w}_{i}$ are weights determined by the displacements $\mathbf{t}_{i}-\mathbf{s}_{i}$, $\phi(\cdot)$ is the radial basis kernel, and $\mathbf{P}(\mathbf{x})$ is a polynomial term for global affine adjustment. We adopt the thin-plate spline (TPS) kernel, defined as $\phi(r)=r^{2}\ln(r)$. In implementation we set $\phi(0)=0$ by continuity, and the previously shown multiquadric form $\sqrt{r^{2}+\epsilon^{2}}$ is corrected here for consistency with TPS.
@@ IV-B Experimental Setup
All experiments were performed on a workstation with Intel Core i9-10900K (3.7 GHz), 32 GB DDR4 memory, and an NVIDIA RTX 3080 GPU (12 GB VRAM). Implementation was done in Python 3.9 using Anaconda environment. Core libraries included NumPy, SciPy, Open3D, scikit-learn, Matplotlib, and OpenCV 4.5. Deep learning inference for semantic segmentation was based on PyTorch 1.13 with CUDA 11.6 support. Algorithm parameters were chosen as follows: keypoint matching threshold = 8 pixels, skeleton extraction density = 40, RBF interpolation parameter $\epsilon=5$, and correspondence distance threshold = 30 pixels. Point cloud preprocessing used 20 nearest neighbors and a std multiplier of 2.0 for outlier removal.
@@ V Conclusion
We presented a post-hoc geo-registration method for terrestrial LiDAR point clouds without reliable GNSS, enabling alignment with satellite imagery. The method also corrects global biases in GNSS-referenced datasets, as demonstrated on KITTI. Through a structured pipeline, our approach aligns road structures between point clouds and satellite images, and integrates elevation data to correct vertical inconsistencies. The method is particularly suited for city-scale geo-registration tasks, where the global topological characteristics of road networks are sufficiently distinctive to resolve matching ambiguities arising from locally repetitive street patterns. The proposed evaluation metric based on the satellite road network provides a generalized and application-oriented means for assessing absolute geo-registration quality at urban scale. Finally, we provide geo-referencing data for the publicly available complete Perth CBD dataset and the five longest sequences of the KITTI dataset to facilitate future research in this direction.